Hello, I have a hard time trying to determine how to exactly go about solving this combination word problem (Please provide an in-depth or even pedantic explanation, as this concept is very confusing for me, thank you):
A play has two male roles, two female roles, and 2 roles that can be either gender. Only a man can be assigned to a male role, and only a woman can be assigned to a female role. If five men and six women audition, in how many ways can the six roles be assigned?

Answer by ikleyn(53765)

.
A play has two male roles, two female roles, and 2 roles that can be either gender.
Only a man can be assigned to a male role, and only a woman can be assigned to a female role.
If five men and six women audition, in how many ways can the six roles be assigned?
~~~~~~~~~~~~~~~~


The roles are M1, M2 (played by men, only);  F1, F2 (played by women, only), and A1, A2  (played by any gender).


There are 5 choices for role M1;  4 choices for role M2; and after that 5-2 = 3 men are in reserve.


There are 6 choices for role F1;  5 choices for role F2;  and after that 6-2 = 4 women are in reserve.


So, now we have 3+4 = 7 persons in the reserve for roles A1 and A2.


The total number of possible combinations is  (5*4) * (6*5) * (7*6) = 25200.      ANSWER


--------------

Solved and carefully explained  ( but not pedantic ).


"Pedantic"  always means  "boring".


Notice that,  although I used the term  "combinations"  in my post,
the order inside these combinations  DOES  MATTER  ( ! )

So actually,  these  "combinations"  are,  in fact,  "permutations"  with some imposed restrictions.


////////////


Do not forget to post your  "THANKS"  to me for my teaching.